# Generative Hierarchical Temporal Transformer for Hand Action Recognition and Motion Prediction

### v2

@@ 4.3 Joint Modeling of Recognition and Prediction
4.4 Modeling Semantic-Temporal Hierarchy
Ours(P.a, w/ $\mathbf{P}$) Ours (P.b, w/ $\mathbf{P},\mathbf{A}$)
@@ 4.4 Modeling Semantic-Temporal Hierarchy
Table 3: Comparison of long-term motion prediction decoded from $\mathbf{m}^{P}$ (P.a) and $\mathbf{m}^{A}$ (P.b), on action sequences that are longer than 2 sec. APD in mm for (left,right) hands.
Action for prediction
We first compare the long-term prediction decoded from $\mathbf{m}^{P}$ and $\mathbf{m}^{A}$, i.e., the two strategies P.a, P.b described in Sec. 3.3, to examine the effectiveness of involving $\mathbf{A}$ for long-term prediction.
As shown in Tab. 3 and Fig. 4, generations from $\mathbf{m}^{A}$ are more realistic and plausible, with lower FID and more consistent global motion. In contrast, results of P.a show larger diversity due to its short-term modeling, but lack fidelity or regularity for long-term motion. Overall, the comparison shows the importance of action modeling in generating faithful and action-guided motions.
Mid-level for prediction
In addition to enabling a decoupled training strategy for $\mathbf{P}$ and $\mathbf{A}$ (Sec. 3.4), the modeling of mid-level features should enhance learning of generation. To verify it, we construct a flattened baseline (i.e., w/o mid-level) by removing the mid-level representation and instead using a single transformer VAE to directly model pose and action. In particular, the flattened baseline has its encoder taking hand poses and object as input and producing the action recognition as output, and has its decoder predicting the future hand poses directly. The flattened baseline has comparable amount of parameters as $\mathbf{P}$ and $\mathbf{A}$, and is trained on the same dataset as our framework for fair comparison.
From Tab. 4 we find that under a comparable FID, the mid-level representation enables better diversity (3rd, 5th cols); meanwhile, as we increase the generation diversity of the flattened baseline via more noisy sampling, its accuracy significantly degrades (4th, 5th cols). From Fig. 4, we can see that the flattened baseline results lack global regularity, despite its modeling of pose and action through a powerful end-to-end transformer VAE. The comparison shows that the mid-level representation enables easier learning of global motion regularity, probably because it decouples the complex task of action-guided motion generation into hierarchical subtasks better captured by $\mathbf{P}$ and $\mathbf{A}$ respectively.
w/o Mid-Level
w/o Mid-Level
Ours (P.b, w/ $\mathbf{P},\mathbf{A}$)
Sampling with $\mu,5\sigma$ $\mu,10\sigma$ $\mu,5\sigma$
@@
Table 4: Comparison of motion prediction, between ours and the ablated setup without modeling the mid-level, on action sequences that are longer than 2 sec. APD in mm for (left,right) hands.
4.5 Discussion
We make more observations about motion prediction, training strategy and mid-level representation, to give additional understanding of the framework. Due to space limit, we provide more details in supplementary.
Training with Assembly101 [42]
We find that including the large-scale Assembly101 for training G-HTT significantly benefits motion prediction on H2O [22] and action recognition, although it is believed that the pose annotation of Assembly101 is not sufficiently accurate for training pose estimators (cf. [35]). The finding points to the importance of large-scale pretraining of fundamental modules.
Predicting action transition
We observe that the model can generate smooth transitions between actions (e.g. motion depicting place spray $\to$ apply spray), although this is never explicitly trained for. It probably comes from training $\mathbf{P}$ on sequences of mixed action annotations (Sec. 3.4), which allows $\mathbf{P}$ to drive the transition by decoding local motions into new actions. Meanwhile, it is also facilitated by the decoupled training of $\mathbf{P}$ and $\mathbf{A}$.
The mid-level regularity
We blend mid-level features of two different input pose sequences, and decode the blended mid-level features into pose sequences. The generated motions naturally interpolate between the tendencies of two given inputs, showing the regularity of the learned mid-level representation.
5 Conclusion
We present a framework to jointly model both recognition and prediction, and leverage the hierarchy of semantic dependency and temporal granularity, for understanding hand pose and action. The framework addresses tasks of input hand pose refinement, action recognition, and future motion prediction, showing improved performances than isolated solutions. To implement the hierarchy, we cascade two transformer VAE blocks, each modeling pose and action respectively, and introduce a mid-level clip-wise motion representation to bridge the two blocks. Each of the pose and action blocks has its encoder and decoder respectively outputting for the recognition and prediction tasks, while the connected cascade enables regular action and motion modeling over both short and long time spans. We train the two blocks separately to flexibly use multiple datasets with different setups and annotation qualities. Extensive tests validate the performance and design of our framework on both recognition and prediction across different datasets.
Limitations and future work
We assume a fixed camera viewpoint for input videos; to process cases with drastic camera movement (e.g., egocentric views with large head motions), an explicit decomposition of hand and camera motion would be necessary, which we leave as future work. Another aspect is to leverage hand motion as priors for robust recognition of manipulated objects, therefore further benefiting action understanding.

### v3

@@ 0.C.2.2 APD
Action Acc.$\uparrow$ - 78.69 69.67
@@
Action Acc.$\uparrow$ - 6.56 77.87
(b) H2O-Test
Resnet-18($\mathbf{F}$) HTT [58] Ours
@@
Action Acc.$\uparrow$ - 85.12 59.92
@@
Action Acc.$\uparrow$ - 12.40 57.85
[…]
Action Acc.$\uparrow$ - 2.89 57.85
(c) AssemblyHands-Val
Resnet-18($\mathbf{F}$) HTT [58] Ours Resnet-18($\mathbf{F}$) HTT [58] Ours
[…]
Action Acc.$\uparrow$ - 16.55 36.01 Action Acc.$\uparrow$ - 9.00 35.52
[…]
Action Acc.$\uparrow$ - 39.42 34.79 Action Acc.$\uparrow$ - 19.95 36.98
[…]
Action Acc.$\uparrow$ - 12.90 33.58 Action Acc.$\uparrow$ - 9.98 36.74
0.C.4 Implementation of PoseGPT [31] (Sec. 4.3 of Main Text)
We adopt the official code of PoseGPT, and make the following modifications for fair comparison in terms of dataset and encoding for pose and action:
  • We train the network by combining the training sets of H2O, Assembly101, and AssemblyHands into one set.
  • We adopt our $\mathbf{H}$ for frame-wise hand pose representation, except that we obtain $p^{L},p^{R}$ by aligning only the root position with the predefined template [43], and therefore having $v^{L},v^{R}\in\mathbb{R}^{3}$ represent the relative translation. The reasons for this root-aligned encoding are that: (1) we observe that our original encoding makes the prediction of PoseGPT suffering from severe jittering; (2) we notice that PoseGPT represents local body-pose via root-aligned global rotations (cf. PoseGPT [31, Sec. 3.1]), which is analogous to our root aligned hand-pose representation.
  • We represent the prescribed action as its CLIP [6, 41, 44] text embedding.
We first train the VQ-VAE $E,D$ for 1200 epochs until convergence, based on which we train the generative model $G$ and observe saturated performance after 30 epochs.
Appendix 0.D Additional Results for Recognition
0.D.1 Complete Version for Tab. 1 of Main Text
To supplement the discussion in Secs. 4.3 and 1 of the main text, we report in Tab. 6 the complete results on all camera views in evaluation, for the H2O-Val, H2O-Test and AssemblyHands-Val splits. The conclusions follow that of the main text, i.e., our results are robustly accurate across different views and datasets.
0.D.2 Additional Qualitative Results for Hand Pose Refinement (Fig. 3 of Main Text)
We provide more qualitative comparison with HTT [58] and image-based estimator $\mathbf{F}$ in the supplementary video. Results show that our G-HTT mitigates jittering present in the input hand motion, and achieves robust generalization across different datasets and camera views.
Appendix 0.E Additional Qualitative Results of Motion Prediction
We provide qualitative cases in the supplementary video, to examine motion prediction from the following aspects:
  • Comparison with baseline methods, adding to Fig. 4 of main text.
  • Prediction diversity. Our model generates diverse motions for a specified action.
  • Prediction from long-term observation. Given a relatively long input sequence, our results show globally action-aware and locally consistent motion.
  • Prediction from image-based estimation. Directly processing a video input, our results again show action-aware and consistent motion.
  • Prediction from Assembly101 annotation. The noisy Assembly101 pose annotations can still be recognized by our model and extended into plausible future motion.
Appendix 0.F Discussion (Sec. 4.5 of Main Text)
0.F.1 Training with Assembly101
We construct the ablated setting by training G-HTT only on H2O and AssemblyHands, excluding the much larger and yet lower-quality Assembly101. We report the comparison of action recognition and motion prediction respectively in Tab. 7 and Tab. 8. Although the ablated setting predicts more plausible motion on the AssemblyHands, by including also the large-scale Assembly101, the network shows enhanced robustness and performance generalization: it outperforms the ablated setting by a significant margin for motion prediction on H2O, and for action recognition on both datasets.
0.F.2 Predicting Action Transition
In the supplementary video, we provide two examples of our predicted motion, which depicts the action transition from place spray to apply spray, and that from grab cappuccino to take out cappuccino.
Table 7: Comparison of action recognition, between ours and the ablated setup trained without leveraging Assembly101.
(a) H2O-Val
Action Acc. cam0 cam1 cam2 cam3 cam4
Training w/o Assembly101
60.66 61.48 63.11 57.38 76.23
Ours 63.11 61.48 69.67 66.36 77.87
(b) H2O-Test
Action Acc. cam0 cam1 cam2 cam3 cam4
Training w/o Assembly101
55.79 58.26 61.16 55.79 57.85
Ours 59.92 65.70 68.18 57.85 57.85
(c) AssemblyHands-Val
Action Acc. v1 v2 v3 v4 v6 v8
Training w/o Assembly101
28.47 28.22 27.25 27.74 26.76 27.01
Ours 36.01 35.52 34.79 36.98 33.58 36.74
@@ 0.F.2 Predicting Action Transition
Table 8: Comparison of motion prediction, between ours and the ablated setup trained without leveraging Assembly101, on action sequences that are longer than 2 sec. Both setups refer to path P.b for motion prediction. APD in mm for (left,right) hands.
H2O-val H2O-Test AssemblyHands-Val
[…]
Training w/o Assembly101 10.72 20.4,25.4 16.29 16.3,34.2 2.69 24.1,36.3
Ours 9.30 27.8,31.4 13.10 24.6,43.3 5.21 33.6,40.0
0.F.3 Predicting from Blended Input
Given two hand pose sequences $\tilde{\mathbf{H}}^{a}_{1:t}$ and $\tilde{\mathbf{H}}^{b}_{1:t}$ with $t=16$ frames, we derive their mid-level feature $\mathbf{m}^{P}_{a},\mathbf{m}^{P}_{b}$, which are the $\mu^{P}$ output by $\mathbf{E}^{P}$. We then obtain $\mathbf{m}^{P}=\lambda\mathbf{m}^{P}_{a}+(1-\lambda)\mathbf{m}^{P}_{b}$, and observe the output motion decoded from $\mathbf{m}^{P}$.
We show an example in the supplementary video, where the two input sequences $\tilde{\mathbf{H}}^{a}_{1:t}$ and $\tilde{\mathbf{H}}^{b}_{1:t}$ show tendencies of screwing and shifting right hand, respectively. By varying $\lambda$ from 1 to 0, the generated motion naturally interpolates between the tendencies of $\tilde{\mathbf{H}}^{a}_{1:t}$ and $\tilde{\mathbf{H}}^{b}_{1:t}$, demonstrating the regularity of the learned mid-level representation.